### v2

# Pulsating with discrete symmetry

Abstract
We propose a class of diffusive lattice models where the internal states of particles are subject to a periodic drive. The discrete symmetry of states enforces an effective energy landscape which can counteract the drive to arrest the dynamics. Such a mechanism is reminiscent of the arrest reported in dense assemblies of pulsating deformable particles, in line with experiments in confluent biological tissues. We show that the competition between arrest and synchronisation promotes an instability which leads to spiral waves. Using analytical coarse-graining, we rationalize the emergence of such spirals in terms of a discrete gauge invariance at the hydrodynamic level.
Active matter describes non-equilibrium systems where a constant energy injection at the particle level leads to collective dynamics far from equilibrium, as is the case for many chemical, biological, and robotic systems [1, 2, 3, 4]. In the last decades, most studies of active matter have focused on the role of self-propulsion, i.e. the ability of each particle to independently undergo directed motion. This paradigm has led to the theoretical understanding of several kinds of collective dynamics which have no counterpart in equilibrium systems [5, 6].
Interestingly, the energy injection in complex units need not result only into a self-propulsion mechanism. An important example is the case of deformable particles [7]. Indeed, complex aggregates such as macromolecules or living cells can change their shape due to internal activity [8], which leads to the spontaneous propagation of contraction waves in dense tissues. Such a wave propagation plays a crucial role in morphogenesis [9, 10, 11, 12, 13, 14], uterine contraction [15, 16], and cardiac arrhytmogenesis [17, 18].
Vertex models are popular to capture the behavior of dense active systems such as biological tissues [19]. They typically consider self-propulsion as the only active component [20, 21]. Yet, when dense tissues behave like solids, it is questionable whether self-propulsion should be the key ingredient. Other models have considered dense assemblies of active deformable particles [22, 23, 24, 25, 26, 27], where energy injection here occurs through the sustained oscillation of size at the single-particle level. With a Kuramoto-like synchronisation [28, 29] of particle sizes, contraction waves spontaneously emerge [26, 27], which are reminiscent of those reported in biological systems [13, 17].
Contraction waves in deformable active particles stem from the competition between synchronisation and steric repulsion [27]. The former favors a global cycling of particle sizes, while the latter promotes an arrested state without any cycling. Combining both stabilizes contraction waves. Interestingly, there are alternative mechanisms for arrest and synchronisation beyond steric repulsion and Kuramoto-like interactions. It remains to explore whether these alternatives lead to a phenomenology similar to that of deformable active particles. Indeed, one may wonder whether the arrest-synchronisation competition actually is a generic scenario for pattern formation. If so, it encourages one to search for any hydrodynamic invariance and/or broken symmetry which might stand out as a hallmark of this competition.
Figure 1: (a) Particles freely diffuse without any volume exclusion. (b) Each particle is subject to an independent periodic driving which pulsates its internal state. Fully-connected onsite interactions locally synchronise the pulsation of particles. (c-d) Instabilities promote the coexistence of pulsating domains. Rotating defects appear at the meeting point of domain interfaces, thus forming spiral waves with a threefold symmetry. $(L,f,\varepsilon,D_{0},\rho_{0})=(512,1.25,2.5,10^{3},10)$.
In this Letter, we formulate a class of pulsating lattice models (PLM) featuring spiral waves with rotating defects [Fig. 1]. We reveal that the discrete symmetry of available states is key to arrest the dynamics, which, in combination with synchronisation and drive, generically yields wave propagation. Through analytical coarse-graining, we map PLM into a specific form of the complex Ginzburg-Landau equation (CGLE) [30] which breaks the continuous gauge invariance, while maintaining a discrete gauge invariance. We demonstrate how such an invariance not only signals the emergence of an arrested phase, but also constrains the features of the waves, at both the hydrodynamic and the particle-based levels. Overall, the arrest-synchronisation competition in PLM provides a generic route for pattern formation, distinct from the well-studied instabilities in other models of reaction-diffusion systems (RDS) [31, 32, 33].
Pulsating lattice models.—We consider $N$ particles in $V=L^{d}$ sites of a hypercubic $d$-dimensional lattice, with global number density $\rho_{0}=N/V$, without any excluded volume. Each site contains an arbitrarily large number of particles. Each particle has an internal state, labeled by a discrete index $a\in\{1,\ldots,q\}$, as a proxy to mimic the internal phase of deformable particles [22, 23, 24, 25, 26, 27]. The crucial difference is that such states now feature a discrete symmetry. In what follows, we focus on the case $q=3$, which is the minimal number of states to accommodate a current, and $d=2$. The system configuration $\{n_{{\bf j},a}\}$ is then given by the number of particles for each state $a$ and site ${\bf j}$.
At every time step $dt/N$, a particle with state $a$ can either jump to a neighbouring site with probability $D_{0}\,dt$, or switch to state $b$ with probability $W_{ba}\,dt$ [Figs. 1(a,b)]:
$W_{ba}=\exp\left[-f_{ba}+\frac{\varepsilon}{\rho_{\bf j}}\left(n_{{\bf j},b}-n_{{\bf j},a}\right)\right].$ (1)
The first term in the exponent of Eq. (1) is defined by $f_{ab}=\pm f$ when $a-b=\pm 1\mod 3$, and $f_{ab}=0$ otherwise. This drive mimics the pulsation of deformable active particles [22, 23, 24, 25, 26, 27]. Inspired by some active lattice dynamics [34, 35, 36, 37, 38, 39, 40], the second term accounts for a synchronising Potts-like [41] interaction with strength $\varepsilon$, where $\rho_{\bf j}=\sum_{a}n_{{\bf j},a}$ is the local density at site $\bf j$. This term favors transition towards the state with the highest number of particles locally.
While interactions between particles are fully connected onsite, different sites exchange information only via diffusion. At small $D_{0}$, many transitions of internal states occur in between two rare jumps of sites. At large $D_{0}$, jumps are so frequent that now all particles are effectively interacting in between two state transitions. In both cases, patterns cannot emerge, and the system can be described solely in terms of the onsite dynamics [42, 43].
Synchronisation and arrest in the onsite case.—In the absence of diffusion ($D_{0}=0$), the evolution of the onsite occupation numbers $n_{a}$ can been studied in terms of the complex parameter
$A(t)=\frac{1}{\rho_{0}}\sum_{a=1}^{3}e^{\frac{2\pi i}{3}a}n_{a}(t).$ (2)
There are three stable phases [42, 43]: (i) a disordered phase at small $\varepsilon$, where particles are uniformly distributed in the three states, with symmetric fixed point $n_{a}=\rho_{0}/3$, and $|A|$ vanishes at large $\rho_{0}$; (ii) a cycling phase at intermediate $\varepsilon$, where particles collectively undergo periodic transitions between states; (iii) an arrested phase at large $\varepsilon$, with three asymmetric fixed points $n_{a}\gg\rho_{0}/3$ invariant under the cyclic permutations $A\to A\,e^{\frac{2\pi i}{3}k}$ ($k$ integer). Transitions between states can here be regarded as unimolecular reactions, in contrast with some models of multimolecular RDS [44, 45]. Besides, the total number of particles here is conserved. These features lead to stabilising disorder, and ensure that the arrest is not an absorbing phase.
In the cycling phase, the discrete nature of the internal states enforces that the global phase $\arg(A)$ is subject to an effective energy landscape. Indeed, the local extrema of the Potts-like interaction, at $\arg(A)=k\pi/3$, constrain the dynamics of the phase. While Kuramoto oscillators (with continuous symmetry) synchronously cycle without any cost, the collective transition of Potts oscillators (with discrete symmetry) between two nearby synchronised configurations implies crossing a desynchronised configuration, which is not energetically favored. Consequently, as $\varepsilon$ increases, Potts oscillators spend more time in a given state before cycling to the next one. Above a critical $\varepsilon$, this energetic effect completely counteracts the drive, breaking down the periodicity of oscillations, and eventually stabilizing the arrested phase [46]. A similar phase trapping has been reported in clock models of aligning self-propelled particles [40].
In short, arrest emerges in PLM solely due to the discrete symmetry of internal states. Therefore, PLM entail a competition between arrest and synchronisation despite the absence of any volume exclusion, in contrast with models of continuously deformable particles [26, 27]. This competition opens the door to the emergence of dynamical patterns in spatially extended systems.
Spiral waves in the spatially extended case.—In the presence of diffusion ($D_{0}>0$), the displacement of particles follows a free dynamics, independently of any interaction, so that the density profile is always homogeneous. Yet, the spatial distribution of the particle states may not remain homogeneous. Indeed, even when particles are synchronised on site, they might not be synchronised between sites, so that the system can potentially accommodate spatial instabilities.
To study the emergence of dynamical patterns, we introduce the local complex parameter
$A_{\bf j}(t)=\frac{1}{\rho_{\bf j}(t)}\sum_{a=1}^{3}e^{\frac{2\pi i}{3}a}n_{{\bf j},a}(t)\equiv r_{\bf j}(t)\,e^{i\psi_{\bf j}(t)}.$ (3)
When all particles are cycling in synchrony, the amplitude $r_{\bf j}\approx 1$ and the phase $\psi_{\bf j}$ are homogeneous. The period of $\psi_{\bf j}$ increases with $\varepsilon$, and eventually diverges. Before diverging, it undergoes large temporal fluctuations, which may desynchronise nearby sites, thus promoting spatial fluctuations of $\psi_{\bf j}$. At large $D_{0}$, spatial fluctuations are suppressed by the rapid displacement of particles in the system. Instead, at moderate $D_{0}$, such fluctuations can potentially build up into a large-scale instability.
Figure 2: (a) Snapshots of the local phase $\psi_{j}$ [Eq. (3), color code in Fig. 1(d)]. (b) Trajectories of the synchronisation parameter $R(t)$ [Eq. (4)]. (c) The averaged parameter $\langle\bar{R}\rangle$ and (d) its standard deviation $\sigma_{R}$ [Eq. (5)] lead to guidelines delineating boundaries between four phases. The solid and dashed lines correspond respectively to $\langle\bar{R}\rangle=0.2$ and $\sigma_{R}=0.045$. Symbols refer to the parameter values taken for panels (a-b) and Fig. 3. $(L,\rho_{0})=(128,10)$, and $f=\varepsilon/2$.
Instabilities can lead to the spatial coexistence of three cycling domains where $r_{\bf j}\equiv r_{{\bf j},c}\approx 1$. Interfaces between domains have a finite width where $r_{{\bf j},c}>r_{\bf j}>0$. The meeting points of interfaces are given by defects where $r_{\bf j}\approx 0$. Since all domains cycle at the same frequency, defects effectively rotate, thus forming spiral waves with a threefold symmetry [Figs. 1(c-d)]. Defects connected by the same interface rotate in opposite directions, and can annihilate by pair when colliding. Higher $D_{0}$ increases the domain sizes and interface widths, thus reducing the number of defects. Higher $\varepsilon$ reduces the interface widths and increases $r_{{\bf j},c}$ [Fig. 2(a)].
Remarkably, we do not observe any phase with turbulent waves, in contrast with the case of continuously deformable particles [27]. Such a turbulence stems in [27] from excitations of the homogeneous arrested phase which promote some localized, aperiodic cyclings. The discrete symmetry of PLM, which entails three arrested states, prevents such events by trapping the phase before it completes one cycle. Therefore, waves in PLM spontaneously organize into steady spirals. Note that the merging of defects can actually also stabilize planar waves. Besides, when initially ordered, the system can also accommodate circular waves without any defect [46].
Figure 3: Probability density of the synchronisation parameter $R$ [Eq. (4)] across the phase boundaries in Fig. 2. (a) The arrested-wave transition and (b) the cycling-wave transition both display a regime of metastability. Distribution measured over 1024 realizations.
Phase boundaries and transitions.—In addition to the phase with waves, we also observe the emergence of three homogeneous phases (disorder, cycles, arrest) analogous to the onsite case. To quantitatively distinguish these phases, we introduce the synchronisation parameter
$R(t)=\frac{1}{N}\bigg{|}\sum_{{\bf j},a}e^{\frac{2\pi i}{3}a}\,n_{{\bf j},a}(t)\bigg{|}.$ (4)
The trajectories of $R(t)$ allow us to identify four scenarios [Figs. 2(a,b)]: (i) a disordered phase at small $\varepsilon$ [$R(t)\approx 0$, yellow line]; (ii) a cycling phase at intermediate $\varepsilon$ and large $D_{0}$ [oscillating $R(t)$, green line]; (iii) an arrested phase at large $\varepsilon$ and large $D_{0}$ [$R(t)\lesssim 1$, blue line]; (iv) spiral waves at large $\varepsilon$ and intermediate $D_{0}$ [$R(t)$ strongly fluctuates, pink line].
To delineate phase boundaries, we consider the time-averaged $\bar{R}$ and the variance $\sigma^{2}_{R}$ of $R(t)$ [Eq. (4)]:
$\bar{R}=\int^{t_{0}+t}_{t_{0}}\frac{R(u)\,\mathrm{d}u}{t},\;\sigma^{2}_{R}=\int^{t_{0}+t}_{t_{0}}\big{\langle}\big{[}R(u)-\bar{R}\big{]}^{2}\big{\rangle}\frac{\mathrm{d}u}{t},$ (5)
[…]
We examine how the distribution $P(R)$ varies across transitions [Fig. 3]. Deep in the wave and arrested phases [resp. pink and dark blue lines], $P(R)$ has a single peak, respectively at small and large $R$. Deep in the cycling phase [green line], $P(R)$ is nonzero for a finite domain of $R$ and peaks at its boundaries, due to the oscillations of $R(t)$. Going from waves to arrest, $P(R)$ becomes bimodal, thus signalling metastability [Fig. 3(a)]. Similarly, going from waves to cycling, $P(R)$ is now nonzero in two separated domains, due to the coexistence between the two dynamical states [Fig. 3(b)]. In short, both the arrest-wave and cycling-wave transitions feature a metastable regime.
Gauge invariance in the hydrodynamics.–Although spiral waves have been already reported in many RDS, one may wonder how the patterns in Figs. 1(c-d) actually differ from standard instabilities present, for instance, in the CGLE [30]. To address this question, we coarse-grain PLM in terms of $A_{\bf j}(t)\to{\cal A}({\bf x},t)$ in the continuum limit. Neglecting density fluctuations, and expanding to lowest orders in ${\cal A}$, we get
$\partial_{t}{\cal A}$ $=D\nabla^{2}{\cal A}+{\cal L}({\cal A})+\Lambda,$ (6)
${\cal L}({\cal A})$ $=c_{1}{\cal A}+c_{2}{{\cal A}^{*}}^{2}+c_{3}|{\cal A}|^{2}{\cal A},$
where $*{}^{*}$ refers to complex conjugation, $\Lambda$ is a zero-mean Gaussian white noise, $D\propto D_{0}$ is the macroscopic diffusion coefficient, and $(c_{1},c_{2},c_{3})$ are complex coefficients fixed by $(f,\varepsilon)$ [46]. The ${{\cal A}^{*}}^{2}$ term in Eq. (6) breaks the continuous gauge ${\cal A}\to{\cal A}\,e^{i\phi}$, in contrast with the standard CGLE [30]. Yet, this term preserves the discrete gauge ${\cal A}\to{\cal A}\,e^{\frac{2\pi i}{3}k}$, in contrast with the hydrodynamics of continuously deformable particles [27]. To the third order in ${\cal A}$, Eq. (6) actually contains all the terms compatible with this gauge.
The local hydrodynamics $\partial_{t}{\cal A}={\cal L}({\cal A})$ reproduces the three homogeneous phases of PLM (disorder, cycles, and arrest) stemming from the interplay between the active drive and the energy landscape [Figs. 4(a-c)]. The breakdown of the continuous gauge ensures the existence of the arrested phase, as in [27]. The discrete gauge enforces that the local hydrodynamics can actually arrest in either one of three states, at variance with [27]. This degeneracy directly affects the shape of the hydrodynamic patterns. Indeed, Eq. (6) entails an instability promoting the spatial coexistence of cycling domains. Rotating defects with threefold symmetry spontaneously form where interfaces meet [Figs. 4(d-e)], yielding the same spiral waves as in PLM [Figs. 1(c-d)].
In short, our coarse-graining shows that PLM are not captured by the standard CGLE [30], and therefore differ from standard RDS. It is striking that, although our hydrodynamics is an expansion at small $\cal A$, the lowest-order nonlinearities are here sufficient to reproduce the specific shape of particle-based spirals. Through the discrete gauge invariance, these nonlinearities bear the hallmark of the microscopic discrete symmetry.
Figure 4: Stream plot of the local hydrodynamics $\partial_{t}{\cal A}=\mathcal{L}({\cal A})$ [Eq. (6)] for (a) disorder, (b) cycles, and (c) arrest. Symbols indicate stable fixed points (green circles), unstable fixed points (red circles), and saddles (yellow squares). (d-e) Spiral waves centered around rotating defects spontaneously form in the hydrodynamics, analogously to the patterns in PLM [Figs. 1(c-d)]. $(L,f,\varepsilon,D,\rho_{0})=(512,0.5,2.5,100,10)$.
Discussion.—We reveal that PLM entail spiral waves stemming from the competition between arrest and synchronisation. Therefore, our results show that the arrest-synchronisation scenario for pattern formation [27] extends to a broad class of models with discrete symmetry. Such a symmetry suffices to promote arrest, thus providing a mechanism distinct from that of repulsive deformable particles [27]. Waves are shaped by the discrete gauge invariance at the hydrodynamic level, which distinguishes them from the patterns of other RDS [31, 32, 33].
Our work paves the way to examining the interplay between discrete symmetry and pattern formation for arbitrary $q$. Indeed, we expect that the three homogeneous phases (disorder, cycles, arrest) are robust beyond $q=3$. To study patterns, one can use a top-down approach postulating the hydrodynamics by identifying the terms which obey the discrete gauge ${\cal A}\to{\cal A}e^{\frac{2\pi i}{q}k}$. As in our coarse-graining [46], one can also derive the hydrodynamic coefficients in terms of the microscopic parameters. It is tempting to speculate that, depending on the parity of $q$, such a hydrodynamic study could lead to identifying generic properties of defect dynamics, and potential connections with the topology of other active models [47].
The discrete symmetry of PLM seems to preclude defect turbulence, at variance with the standard CGLE [30] and its recent generalization [27]. To capture such a turbulence, one could introduce energy levels which maintain the discrete nature of PLM while breaking its symmetry. The arrested phase would no longer be degenerate, thus opening the door to local excitations nucleating defects from a homogeneous phase, as reported in [27]. In this context, it would be interesting to explore how density fluctuations affect defect nucleation, at both the hydrodynamic and the particle-based levels.
Acknowledgements.
We acknowledge fruitful discussions with L. K. Davis, M. Esposito, J. Meihbom, and Y. Zhang. This project has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101056825 (NewGenActive), and from the Luxembourg National Research Fund (FNR), grant reference 14389168.